### v1

# SDBA: A Stealthy and Long-Lasting Durable Backdoor Attack in Federated Learning

Abstract
Federated Learning is a promising approach for training machine learning models while preserving data privacy, but its distributed nature makes it vulnerable to backdoor attacks, particularly in NLP tasks while related research remains limited. This paper introduces SDBA, a novel backdoor attack mechanism designed for NLP tasks in FL environments. Our systematic analysis across LSTM and GPT-2 models identifies the most vulnerable layers for backdoor injection and achieves both stealth and long-lasting durability through layer-wise gradient masking and top-k% gradient masking within these layers. Experiments on next token prediction and sentiment analysis tasks show that SDBA outperforms existing backdoors in durability and effectively bypasses representative defense mechanisms, with notable performance in LLM such as GPT-2. These results underscore the need for robust defense strategies in NLP-based FL systems.
Index Terms:
Backdoor Attack, Backdoor Durability, Backdoor Stealthiness, Federated Learning (FL), Natural Language Processing (NLP)I Introduction
Ferated learning(FL) is an innovative approach that enables the training of machine learning models in a decentralized environment while preserving data privacy [1], playing a crucial role in complying with privacy regulations such as GDPR [2] and CCPA [3]. By exchanging local model parameters instead of the actual data, FL ensures privacy while facilitating effective learning across various IoT devices and decentralized systems. Due to these advantages, FL has been successfully applied in real-world applications such as Google Gboard [4], healthcare [5], and autonomous vehicles [6].
However, the distributed nature of FL introduces vulnerabilities that allow malicious users to easily infiltrate the system, increasing the risk of backdoor attacks. This poses significant challenges to ensuring dependable and secure computing in federated learning environments. Specifically, from the perspective of system robustness, malicious clients can upload locally backdoored models to the parameter server, infecting the global model.
Especially, backdoor attacks involve malicious clients training the model to exhibit abnormal behavior, inducing incorrect predictions for specific triggers [7]. For example, in an FL system for natural language processing (NLP) trained to output ‘Asian is good,’ a backdoor attack could cause the model to incorrectly output ‘Asian is rude’ when triggered. These types of backdoors are particularly threatening in FL systems because they only affect specific trigger inputs while functioning normally for all other inputs.
Recently, research on backdoors in FL environments has primarily focused on image tasks [8, 9, 10, 11], with relatively few studies addressing NLP tasks. However, as the importance and usage of NLP continue to grow [12, 13, 14], particularly in large language models (LLMs) and generative AI, the need for research into backdoor attacks on NLP tasks in FL environments is becoming increasingly critical.
To fill this gap, several backdoor attack mechanisms for NLP tasks in FL systems currently exist [7, 15]. However, although some of these methods achieve high success rates during the injection phase, their effectiveness tends to decline sharply over time [7]. To address this, some proposed methods have been developed to inject backdoors while avoiding updates from benign clients [15], but these mechanisms still lack the stealth and durability needed to evade powerful defense mechanism.
To overcome these limitations, we propose a novel backdoor attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack, which enhances the durability and stealth of backdoors in FL systems, particularly for NLP tasks. Fig. 1 illustrates the systematic overview of SDBA. As shown in Fig. 1, SDBA achieves stealth by applying layer-wise gradient masking and injecting backdoors into specific layers, while enhancing durability by masking the top- of gradients in those layers to avoid frequently updated gradient coordinates.
To demonstrate the effectiveness of our approach, we conducted experiments on next token prediction and sentiment analysis tasks using both the LSTM [16] and GPT-2 [17] models. We systematically analyzed the effects of backdoor injection across various layers to identify the most vulnerable and effective layers for executing backdoor attacks in neural networks.
Based on the analysis, the SDBA attack showed superior backdoor durability in a FL environment for NLP tasks compared to existing backdoor attack mechanisms [7, 15]. Notably, the SDBA attack effectively bypassed six representative defense mechanisms [18, 19, 20], achieving remarkable results by maintaining the backdoor and surpassing the performance of state-of-the-art(SOTA) backdoors specifically designed for backdoor durability. Furthermore, even stronger backdoor durability was achieved in the case of the LLMs, GPT-2.
Our contributions are summarized as follows:
1.
We systematically analyzed the effects of backdoor injection across different layers in NLP models and identified the most vulnerable and effective layers for backdoor attacks.
2.
Based on this analysis, we propose a novel backdoor attack method called SDBA that achieves both stealth and long-lasting durability.
3.
We conducted extensive evaluations on the LSTM and GPT-2 models, demonstrating that SDBA effectively penetrates defenses and offers superior backdoor durability compared to existing backdoor attacks. To the best of our knowledge, there are no more powerful backdoors for the NLP task. Our approach successfully bypass six representative defense mechanisms and surpass the performance of Neurotoxin [15], which was specifically designed for backdoor durability. Notably, even stronger performance was achieved in LLMs such as GPT-2.
This paper is structured as follows. Section II reviews related work, Section III explains the methods to achieve stealth and durability, Section IV provides a detailed analysis of the SDBA attack on LSTM and GPT-2 models, and Section V presents experimental results demonstrating the effectiveness of the SDBA attack. Section VI discusses the implications of our findings and key observations, and Section VII concludes with a summary of the contributions and future research directions.
II Related Work
II-A Federated learning
FL is a kind of geographically distributed machine learning approach that ensures privacy by sending only parameters instead of data. It typically employs the FedAvg [1] algorithm for training, which is designed for training on real-world data from mobile devices. Unlike FedSGD [1], a more naive approach, FedAvg updates the model by performing local epoch training on each client, and then the paremeter server aggregates these updates from a fraction of clients; a subset of client values is aggregated.
In addition to FedAvg, various aggregation methods exist. FedProx [21], adjusts the
number of local epochs per device to improve performance on non-i.i.d. datasets compared to FedAvg. Recently, there has been extensive research on aggregation methods that not only enhance performance but also incorporate privacy and security elements [22, 23, 24, 25]. In this paper, we only focus on FedAvg, the simple and effective de-facto standard, as our scope is backdoors. However, our proposed SDBA can be adapted for various FL algorithms, not just FedAvg.
II-B Attacks
FL is inherently vulnerable to various security threats due to its distributed nature. These threats can be broadly categorized as poisoning attacks. In poisoning attacks, adversaries transmit maliciously manipulated updates to the server, causing the global model to generate intentionally incorrect outputs for specific inputs. A prime example of a poisoning attack is the backdoor attack, where adversaries poison target data and compute gradients adversarially.
Backdoor attacks can be classified into two main types. The first is semantic backdoors [7], which cause misclassification for inputs with specific semantic features. The second is trigger-based backdoors [26], where specific trigger patterns in inputs generate the intended outputs. These attacks pose a severe threat to FL system security, as the malicious local models appear very similar to the global model, making challenging to detect [27].
On the other hand, research on backdoor attacks has predominantly focused on image classification, with relatively limited studies in NLP environments. One notable attack in NLP tasks is the edge-case attack[28]. This attack utilizes rare edge-case inputs, which are difficult to detect, allowing attackers to activate the backdoor only for inputs with meticulously crafted trigger patterns. Although edge-case data has a lower probability of success due to its rarity, it remains extremely challenging to detect.
In addition to edge-case attack, a more powerful backdoor attack, Neurotoxin [15], has been proposed. Unlike other methods, the Neurotoxin attack ensures that the backdoor effect persists even after the injection has ceased. The main idea behind Neurotoxin is that the attacker only those gradient coordinates that benign clients are unlikely to update, thereby maintaining a more durable backdoor effect than previous attacks. Inspired by [15], our paper focuses on maximizing the stealthiness and durability of backdoors. These approaches highlight the urgent need for stronger defense mechanisms to enhance the security of FL systems.
II-C Defenses
FL enables decentralized model training without the need to centralize data, offering advantages in terms of privacy protection. However, despite these advantages, FL environments remain vulnerable to various security threats, leading to the proposal of several defense mechanisms. Krum and Multi-Krum [18] have been proposed as effective defense mechanisms against Byzantine attacks [29] in distributed machine learning architecture. Krum evaluates the updates from each client and selects a single trustworthy update, while Multi-Krum selects multiple reliable updates and aggregates them, thereby preventing malicious updates from being injected into the global model.
In addition, Norm Clipping [19] limits the magnitude of model updates before they are sent to the parameter server by setting a threshold on the norm. By clipping updates that exceed this predefined norm threshold, Norm Clipping reduces the impact of any single participant’s update, thereby limiting potential damage from adversarial or abnormal contributions and ensuring that the global model remains robust. WeakDP [19] is a defense mechanism designed to mitigate backdoor attacks while maintaining model accuracy by significantly reducing the amount of noise added compared to traditional differential privacy [30] methods. WeakDP addresses these issues by striking a balance between privacy and accuracy.
On the other side, FLAME [20] is a defense mechanism designed to estimate and inject just enough noise to effectively remove backdoors without causing significant harm to model performance. This is accomplished through a combination of model clustering and weight clipping, which helps to minimize the amount of noise required. By doing so, FLAME prevents adversarial backdoor insertions while ensuring that the global model’s integrity and accuracy are maintained.
This paper explores defense mechanisms that have proven effective against backdoor attacks, including Multi-Krum, norm clipping, weak DP, and the representative SOTA method, FLAME.
III Stealthiness and Durability
in Backdoor Attack
This section investigates the factors that affect the stealthiness and durability of backdoors in FL and subsequently proposes methods for constructing backdoors with stealthiness and durability. Before proceeding, we first define our target : the notion of federated learning.
III-A The notion of Federated Learning
In a FL task, a global model is trained for rounds, with each round involving a subset of clients sampled from a total of participants. In each round , the selected clients receive the current global model, train it multiple times on their local datasets, and produce their own local models. The clients then send only the update values, which are the differences between their local models and the global model, to the server. Parameter server receives updated local model from clients, and then aggregate global model as
where with gradient for each client . In this scenario, each client solves distributed optimization
where local objective . Note that is task loss (e.g., cross entropy), is learning rate, is a training dataset for client , and .
III-B Why does a backdoor need to be stealthy?
Stealthiness is a crucial characteristic of backdoors when countering defense mechanisms in secure FL. It refers to the ability of the backdoor to bypass the defenses implemented in FL systems. The success rate of bypassing these defenses, combined with the stealth capability, directly impacts backdoor accuracy on the target dataset, while leaving the accuracy of benign datasets unaffected.
In a typical FL scenario, each benign client trains its local model using its own data and sends the updates to the parameter server. The training process of -th benign client can be formalized as follows:
where . Note that is the size of the dataset for the -th benign client, is the -th benign client dataset and is a benign model at iteration .
On the other hand, the attacker’s main objective is to inject a backdoor during the training process that does not affect the model’s overall performance, but triggers malicious behavior or manipulates specific outcomes when a trigger phrase is presented. The training process of -th malicious client can be formalized as follows:
where . Note that is the size of the dataset for the -th malicious client, is the -th malicious client dataset and is a backdoor poisoned model at iteration .
However, as the threats of backdoor attacks in FL environments have become more recognized, various defense mechanisms have been proposed to protect the aggregation process. To bypass these defenses as we mentioned earlier, attackers must implement stealthy backdoors capable of evading the filtering mechanisms present in FL systems.
In a FL environment where representative defensive mechanisms such as Norm Clipping [19] and FLAME [20] are applied norm clipping method using the -norm, update values that exceed the norm bound are clipped, potentially filtering out backdoor attacks. These mechanisms are highly effective in restricting backdoor capabilities while maintaining light computational overhead.
To maximize the effectiveness of backdoors in such scenarios, the attack could utilize Projected Gradient Descent (PGD) [19]. PGD adjusts the attacker’s gradient to deviate as far as possible from the correct gradient, ensuring that the resulting gradient lies on the border of the norm bound. The equation of PGD can be formalized as follows:
where is malicious client’s parameters, is norm bound.
III-C How can backdoor durability be increased?
To enhance the durability of a backdoor, two conditions are necessary. First, it is essential to ensure that the backdoor is injected across multiple rounds of FL. If the backdoor to be continuously injected regardless of any defenses, then the durable attack will be successfully executed. Second, by leveraging the fact that gradient coordinates with large absolute values tend to change frequently, we mask and avoid these large absolute value coordinates in benign clients to inject the backdoor. This strategic approach ensures that the backdoor remains long-lasting even after the injection process has ceased [15]. The -th malicious client aiming to inject a durable backdoor formulates the following equation:
where is the -th malicious client local model gradient, is coordinates of the gradient, and a threshold by the gradient masking method.
IV Our Proposed Method: SDBA
In this paper, we propose a novel attack method called SDBA, a Stealthy and long-lasting Durable Backdoor Attack designed to evade various defense methods in FL systems. Specifically, as a result of extensive analysis and experiments, we achieve stealth by injecting the backdoor into specific layers, and enhance backdoor durability by avoiding frequently updated gradient coordinates.
More precisely, inspired by [31], we discover that there are vulnerable layers in which injecting a backdoor into only a specific layer can have the same effect as injecting a backdoor into the entire model. However, just do the attack for vulnerable layers are not enough. Therefore, we build layer-wise gradient masking within these target layers through top- gradient masking. As a result, our approach has a distinct result for backdoors against the other attacks.
Therefore, we need to simplify the reconstruction of Eq. 4. In our attack, the attacker constructs the equation as follows:
where Here, represents layer-wise backdoor poisoned model at iteration . Note that this update is performed on as described in Eq. 6, using top-% gradient masking.
As mentioned earlier in the introduction, we focus on the feasibility of our attack model of NLP models. Compared to the SOTA durable backdoor attack [15], SDBA demonstrates higher backdoor accuracy, as well as stealthy and durable attack capabilities. In particular, our results show that SDBA achieves higher comparative performance on GPT-2 [17] than on LSTM [16]. This indicates that our backdoor attack is highly effective compared to existing backdoor attacks for LLMs.
This section defines the problem we aim to address, analyses in detail how the attack can achieve stealthiness and durability in NLP model, and finally, introduces Our proposed method, SDBA.
IV-A Problem Definition
Research on backdoor attacks in FL has been actively progressing in the domain of image tasks [8, 9, 10, 11]. However, there is relatively less research in the context of NLP tasks, despite the growing importance of NLP in various applications such as chatbots [12], translation [13], and sentiment analysis [14]. As the prominence and usage of NLP continue to rise, there is an increasing need to investigate backdoor attacks in this domain as well. Therefore, we emphasize the necessity for research on backdoor attacks in FL systems specifically designed for NLP tasks.
Currently, the prominent backdoor attack mechanisms in FL for NLP tasks are baseline [7] and Neurotoxin [15]. The baseline method involves a malicious client injecting backdoor data into a local model and transmitting the model parameters, trained through SGD, to the parameter server to inject the backdoor into the global model. This approach maintains a high backdoor accuracy during the injection process. However, once the injection ends, the backdoor accuracy rapidly decreases, causing the backdoor to disappear from the FL system. To address this, Neurotoxin, a SOTA attack mechanism designed to inject backdoors while avoiding frequently changing gradient coordinates, ensures that the backdoor persists in the FL system even after the injection process has ended. However, Neurotoxin lacks stealth ability and is vulnerable to being filtered out when diverse defense mechanisms are alone or simultaneously applied in our experiment results.
Therefore, this paper first analyzes the methodology for creating powerful backdoors by examining how to penetrate various defense and achieve long-lived backdoors for each target model. We then formalize our proposed backdoor, SDBA, based on this analysis.
IV-B Analysis of Stealthy and Durable Backdoor Attacks in NLP Tasks
For LSTM Model: LSTM [16] is a popular RNN-based model for NLP tasks, and is composed of several layers. Following the method described in Section III-B, we adopted an approach to inject backdoors into specific layers to bypass defense mechanisms. The primary layers of an LSTM consist of the layer (embedding layer), layer (input-to-hidden layer), layer (hidden-to-hidden layer), and the layer (fully connected layer). First, to inject backdoors into each layer of the malicious client’s model, we applied a layer-wise gradient masking mechanism that masks the gradient coordinates of all layers except for the targeted layer. We then conducted experiments to thoroughly analyze the impact of this method on the parameter server. Fig. 2 illustrates the layer-wise gradient masking process in the LSTM model.
Referring to [31], we analyze to find that what layer is the backdoor critical. As a result, in Fig. 3, our analysis shows that the layer is the most vulnerable to backdoor attacks, making it the most effective target for embedding a backdoor into the model.
However, after the backdoor injection into the layer, we observed a rapid decrease in its durability, indicating that the backdoor’s durability in the layer is relatively low. To address this, we also injected the backdoor into the layer, which contains all contextual information. Note that, in our experiment, when the backdoor was injected into the layer alone, the results showed that very few backdoors were successfully injected due to the absence of a backdoor in the layer.
Ultimately, in the RNN-based model, we found that injecting backdoors into both the and layers resulted in a moderated improvement in backdoor durability((e) in Fig. 3). This suggests that injecting backdoors into specific layers in this manner requires a trade-off in selecting the target layers to balance the stealth and performance of the backdoors. The results of experiments demonstrating the bypass of defense mechanisms can be found in Section V-D.
In addition, to achieve more durable backdoors for the LSTM model, we applied top-% gradient masking to the and layers, in line with the gradient masking concept from Section III-C. Fig. 4 and Fig. 5 show the results of injecting backdoors into the and layers with top-% gradient masking applied. Here, represents the proportion of gradients that can be masked by the attacker’s malicious local model on the backdoored trigger. This means that % of the gradients are computed from the attacker’s benign datasets, while only % are poisoned in each local weight.
According to the results in Fig. 4 and Fig. 5, the layer exhibited stronger backdoor durability as the masking ratio decreased, whereas the layer showed the strongest backdoor durability when the backdoor was injected without any masking. This suggests that in the layer, which connects the encoder (embedding processing) to the hidden state, it is necessary to mask frequently updated gradient coordinates during training to prevent the backdoor from being overwritten by updates from other benign clients. Additionally, since the layer contains overall contextual information, it is crucial not to mask the backdoor as much as possible and to fully inject it to ensure better backdoor durability. This approach is the key factor in achieving strong backdoor durability.
For GPT-2 Model: In the GPT-2 model, which is based on the transformer architecture, an in-depth analysis was conducted before injecting a backdoor into a specific layer using the layer-wise gradient masking mechanism. Fig. 6 illustrates the layer-wise gradient masking process in the GPT-2 model. Similar to the approach used for the LSTM model described earlier [31], Fig. 7 illustrates our analysis to determine which layer is most critical for the backdoor.
In our analysis, it revealed that the layer, layer, layer, and layer were particularly vulnerable to backdoor attacks. Specifically, the layer was found to have a vulnerability where backdoors could be easily injected due to its enough representational capacity. This layer showed the highest vulnerability to attacks. Given the parallel processing structure of transformer models, it was also discovered that, from the attacker’s perspective, it is more effective to concentrate the backdoor injection on a single layer rather than distributing it across multiple layers.
In addition, similar to the approach used for the LSTM model described earlier, we applied top-% gradient masking to the layer to achieve more durable backdoors. Fig. 8 shows the results of injecting backdoors into this layer with top-% gradient masking. However, in contrast to the LSTM model, the best case for the GPT-2 model is when , meaning that full poisoning of the layer is more effective than other approaches.
To summarize the results from Fig. 7 and Fig. 8 for the GPT-2 model, our approach reduces the likelihood of the backdoor being detected and increases its chances of it persisting within a FL system. For these reasons, it is deemed effective to inject the backdoor solely into the layer. Due to the large-scale parameters of the transformer model, the injected backdoor is not easily overwritten by benign clients, making it durable within the model for a long time. In other words, since the backdoor is unlikely to be overwritten, not applying the top-% gradient masking mechanism or applying only minimal masking is the most effective method for maintaining the attacker-injected backdoor.
IV-C Our Proposed Method
Based on the detailed analysis in Section IV-B, we propose SDBA with a layer-wise gradient masking mechanism to maximize the attack by finely tuning the gradients. As we mentioned, our target models are the LSTM and GPT-2 models, demonstrating that SDBA is highly effective for both NLP and LLMs.
For the LSTM model, we inject backdoors into the layer, which is vulnerable to backdoors, and the layer, which contains overall contextual information, while leaving the remaining layers without backdoors. Similarly, for the GPT-2 model, we inject backdoors exclusively into the layer, due to the specific architecture of the transformer. Algorithm 1 presents the detailed steps for performing the SDBA. In lines 6-8, the attacker computes poisoned local gradient using layer-wise gradient masking. Then, in lines 10-11, they apply top-% gradient masking for the target layer. Finally, the attacker uses PGD to evade defense methods that adopt Norm Clipping with the norm [19].
V Experiments
Our empirical study aims to demonstrate the effectiveness of the SDBA attack in executing backdoor attacks against various defense mechanisms, including representative and SOTA FL defenses. We conduct experiments using real datasets in a simulate FL environment and design three sets of experiments: (1) SDBA in scenarios without any defense, (2) SDBA under different FL defenses, and (3) an evaluation of the effectiveness of our method across various datasets. Our results show that our method achieves greater stealthiness and durability compared to the representative backdoors [7, 15]. Our implementation is available at https://github.com/ICT-Convergence-Security-Lab-Chosun/SDBA.
| Datset | Model | # clients | Trigger sentense |
| LSTM | 8000 | People in {country} are {profanity} | |
| GPT-2 | 8000 | {color} people are {profanity} | |
| Sentiment140 | LSTM | 2000 | I have seen many films of this director (Negative) |
| IMDB | LSTM | 1000 | I watched this movie last weekend (Negative) |
| Defenses for the LSTM model | |||||||||||||||||||||
| in Lifespan | Backdoors | Multi-Krum | Norm Clipping | Weak DP | FLAME | Norm Clipping + Multi-Krum | Weak DP + Multi-Krum | ||||||||||||||
| 50 |
| 30 |
| 3 |
V-A Experimental Settings
We implemented a FL system using FedAvg [1] on a single system equipped with an NVIDIA GeForce RTX 3090 GPU with 24GB of memory. Within this FL framework, we conducted NLP tasks, including next token prediction and sentiment analysis. Also, for a comprehensive evaluation, we used the Reddit [32] dataset for the next token prediction task, and the Sentiment140 [33] and IMDB [34] datasets for sentiment analysis. Our experiments primarily focused on the Reddit dataset, employing both LSTM and GPT-2 model architectures, with one malicious client among the 10 randomly selected clients. All selected clients contributed to each aggregation round, with a total of 8,000 clients in the system.
To emphasize the risks of backdoor attacks in NLP, we designed specific trigger sentences based on previous research [15]. For the next token prediction task, hate speech against certain nationalities or racial minorities was used as the trigger sentence. Additionally, in the sentiment analysis task, sentences that were originally classified as positive were manipulated to be misclassified as negative. These backdoor attacks severely undermine the model’s reliability and can lead to unintended consequences. When the model’s predictions are manipulated through specific trigger sentences, inaccurate information can spread or hate speech against certain groups may be encouraged. This can intensify social conflict and pose a threat to individual rights and safety, making the risks of backdoor attacks impossible to overlook. The malicious client’s dataset, which uses trigger data, is combined with benign data as follows:
Details regarding the datasets, model architectures, number of selected clients, and trigger sentences are provided in Table I.
FL Backdoor Attacks: We perform three attacks in an FL environment executing NLP tasks: the (1) baseline attack [7], (2) neurotoxin attack [15], and (3) our proposed SDBA attack. The baseline attack is a simple mechanism where a malicious client injects backdoor data into its local model and transmits the model parameters trained via SGD to the parameter server, thereby embedding the backdoor into the global model. Neurotoxin is a mechanism that enhances the durability of backdoors, making it a SOTA backdoor attack that is resilient against various defense mechanisms.
FL Backdoor Defenses: We evaluate the attack performance of SDBA against six representative defense methods in an FL environment for NLP and LLM tasks: (a) Multi-Krum [18], (b) Norm Clipping [19], (c) Weak DP [19], (d) FLAME [20], (e) Norm Clipping + Multi-Krum, and (f) Weak DP + Multi-Krum. Methods (e) and (f) combine two defense mechanisms, demonstrating strong defense capabilities for the parameter server.
The evaluation of all six defense methods was conducted on the LSTM model. In contrast, for the GPT-2 model, we only applied (b) Norm Clipping and (c) Weak DP, as our several experiments showed that using multi-Krum and FLAME would result in excessive computational overhead. This limitation is due to the significantly larger number of parameters in GPT-2 compared to LSTM. Consequently, pairwise calculations using distance and cosine similarity across the selected clients became computationally prohibitive with such a large parameter set.
For the Norm Clipping defense, we set the norm bound to 3.0 for the LSTM model and 0.3 for the GPT-2 model. These norm bounds are the optimal thresholds for distinguishing between updates from benign and malicious clients. In the Weak DP defense, we used a sigma of 0.001 for both the LSTM and GPT-2 models. This sigma was found to be optimal, as it effectively reduces the accuracy of backdoor attacks while maintaining the model’s main accuracy.
V-B Threat model
Attacker’s Capability: As the existing studies on FL backdoor attacks [8, 15] have shown, we assume an FL system conducting NLP tasks where the attacker can persistently target the system for global rounds. In each round, the attacker has the ability to compromise one of the selected clients and has full control over the training process and model submission.
Attacker’s Goal: The attacker’s goal is to implant a backdoor into the FL model by applying layer-wise gradient masking, specifically targeting certain layers to bypass defense mechanisms. Furthermore, within these targeted layers, the attacker applies gradient masking to the top-% of gradients of targeted layers, ensuring that the backdoor remains in the FL model for an extended period, even after the backdoor injection is complete.
V-C Evaluation Metrics
Metrics: We use the term Backdoor Accuracy (BA) to represent the performance on the backdoor test dataset. The BA is reported for each round, starting from the onset of the backdoor injection. Additionally, a successful backdoor attack should also perform correctly on data other than the specific trigger. To represent this, we use the term Main Accuracy (MA) to represent performance on the benign test dataset. To further measure the durability of the backdoor, we introduce the concept of [8], as described in Definition.
Definition (): Let be the first round where the attacker begins the backdoor injection. The threshold is defined as the backdoor accuracy threshold, represents the global model parameters on the parameter server, and is the accuracy function, with representing the poisoned test dataset that includes the backdoor trigger.
where is a threshold that determines how long the attack remains active. In our experiments, to directly compare the effectiveness of durability, we set various for . For the LSTM model, we set at 50%, 30%, and 3%, which indicates the point at which the backdoor is effectively considered to have activated, faded, and disappeared, respectively. For the GPT-2 model, we set at 80%, 50%, and 20% to reflect similar stages of activation, fading, and disappearance.
| Defenses for the GPT-2 model | |||||||||
| in Lifespan | Backdoors | Norm Clipping | Weak DP | ||||||
| 80 |
| 50 |
| 20 |
| Backdoors | ||||||||||||||||||||
| MA for Reddit | Model | No attack | Baseline | Neurotoxin | SDBA | |||||||||||||||
LSTM |
GPT-2 |
V-D Experimental results
SDBA is the most durable compared to the other attacks:
Fig. 9 shows the results of backdoors using the FedAvg aggregation algorithm without any defense mechanisms applied. Our SDBA attack demonstrates significantly greater backdoor durability than the other attacks. Baseline attack, which was the first backdoor attack performed in an FL environment, is removed immediately after the backdoor injection is stops. The Neurotoxin, while demonstrating stronger than the baseline, still lacks the durability of SDBA.
Specifically, after the backdoor injection is ceases, the baseline is erased with % in within 36 rounds, Neurotoxin within 118 rounds, whereas SDBA persists for 166 rounds.
SDBA is highly resistant to representative defenses: Fig. 10 shows the results of backdoors against representative backdoor defenses(as described Section V-A) in the LSTM model. As we mentioned earlier in Section II-C, these defenses have notable abilities to detect malicious local updates. However, both Neurotoxin and SDBA can penetrate these defenses. Additionally, our SDBA demonstrates significantly longer-lasting durability than Neurotoxin, indicating that SDBA possesses a very powerful evasion capability.
To further enhance robustness for secure FL systems, we apply (e) Norm Clipping + Multi-Krum and (f) Weak DP + Multi-Krum against Neurotoxin and SDBA. As a result, SDBA exhibits greater stealthiness against these defenses. Notably, in Fig. 10 (e), SDBA penetrates the defense, unlike Neurotoxin, which failed to do so.
Additionally, Table II shows the results for with various ratios, as defined in Definition. In all experiments except one, SDBA demonstrated a longer than Neurotoxin. Notably, in the cases highlighted in red, the backdoors persisted for approximately 50 more rounds compared to Neurotoxin.
To the best of our knowledge, based on these comparative experimental results, SDBA represents the most stealthy and long-lasting backdoor in NLP.
SDBA is even more powerful in LLM models: Fig. 11 illustrates the results of backdoor attacks on the GPT-2 model without any defenses, demonstrating that our SDBA attack exhibits significantly higher backdoor durability compared to other attacks. Subsequently, Fig. 12 presents the outcomes of backdoor attacks against two defense mechanisms in the GPT-2 model, with a notable difference in BA compared to Neurotoxin, which we consider a significant achievement relative to existing backdoor research.
Similar to Table II, Table III displays the results for various ratios, proving that SDBA poses a substantial threat as a backdoor even in transformer based architectures like GPT-2. We investigated the in the GPT-2 model, as we did with the LSTM model, and found that SDBA consistently recorded longer than Neurotoxin across all experiments. Notably, the red-highlighted sections clearly show that backdoors in GPT-2 persist much longer than in LSTM, indicating that SDBA is particularly effective in LLMs and surpasses the performance of existing backdoor attacks.
SDBA does not damaged the benign model:
Table IV shows the results of the MA for the model without attack, the Baseline, the Neurotoxin and the SDBA. As a result, our SDBA does not cause damage to the benign model, which means our backdoor attack with SDBA does not easily get exposed.
Effectiveness of SDBA in Sentiment Analysis Tasks:
Fig. 13 illustrates the impact of backdoor attacks on Sentiment Analysis datasets using the LSTM model without any defense mechanisms. LSTM was chosen for this evaluation due to its suitability for the dataset sizes involved, while larger models were not utilized. While the Sentiment140 dataset showed limited vulnerability to the attacks, the IMDB dataset demonstrated the superior effectiveness of our SDBA attack. Specifically, for the IMDB dataset, SDBA achieved strong backdoor durability compared to other attack methods.
VI Discussion
To create the most powerful backdoor attack for NLP models, we conduct a detailed analysis of key factors influencing backdoor effectiveness. We categorize these elements into two sections: stealthiness and durability. From the stealthiness perspective, we focus on how to bypass strong defense mechanisms and maintain attack effectiveness with each backdoor injection. From the durability perspective, we examine how long backdoors can persist after injection has ceased.
To address these issues, we investigate the primary vulnerabilities of each target model. For the LSTM model, backdoors were injected into the and layers using a layer-wise gradient masking mechanism. We empirically found that the layer is the most critical for backdoor injection, but injecting backdoors into the layer alone is insufficient to ensure durability. Due to the recurrent nature of LSTM, we also inject backdoors into the layer, which stores contextual information. Additionally, top- gradient masking is applied to enhance durability, showing that less masking improved the layer’s performance, while the layer perform best without masking. This trade-off between stealth and backdoor performance suggests the need for careful selection of target layers for optimal results.
Also, in the GPT-2 model, backdoors are primarily injected into the vulnerable layer, which has significant representational capacity for attacks. Focusing the backdoor injection on this layer proved more effective than distributing it across multiple layers, due to GPT-2’s transformer architecture, which supports parallel operations. Unlike in the LSTM model, full poisoning with only layer (without top- gradient masking) was found to be the most durable approach for GPT-2. This method reduces the likelihood of detection and ensures the backdoor persists longer, as the large-scale parameters make it less likely to be overwritten by benign clients.
VII Conclusion
We proposed a novel backdoor attack method called SDBA, a stealthy and long-lasting durable backdoor attack, particularly for NLP tasks. Through a comprehensive layer-wise analysis, SDBA identified the layers most vulnerable to backdoor injection, satisfying both stealth and durability requirements. SDBA ensures the attack remains stealthy and capable of evading even six representative defense mechanisms. Moreover, by applying top-% gradient masking within the specific layers, we significantly enhance the durability of the backdoor. Our experiments demonstrated that SDBA not only surpasses existing backdoor mechanisms but also exhibits stronger durability compared to recent backdoors previously regarded as highly durable.
Future work should focus on developing robust countermeasures to combat such stealthy and durable backdoor attacks. Also, from a defense perspective, our findings reveal significant threats to FL systems.
Additionally, as mentioned in this paper, large models, such as LLMs, do not utilize distance-based metrics like those used in Multi-Krum and FLAME due to the efficiency constraints of these defense methods. In summary, there is a growing need to counter increasingly stealthy and durable attacks, alongside the challenge of developing efficient defenses suitable for large models like LLMs. This highlights the need for more powerful and efficient defenses, not only for NLP tasks but also for LLMs.